OpenStudy (superhelp101):

When solving x2 + 6x = 10 by the completing the square method, what value is added to both sides? 9 5 4 3

OpenStudy (superhelp101):

@jim_thompson5910

OpenStudy (superhelp101):

@satellite73

OpenStudy (superhelp101):

@Hero @ganeshie8

OpenStudy (superhelp101):

@zzr0ck3r @SithsAndGiggles

OpenStudy (superhelp101):

@dumbcow

OpenStudy (superhelp101):

@aum I need your help

OpenStudy (superhelp101):

@Luigi0210 @phi

OpenStudy (anonymous):

How would you get \((x+3)^2\) on the right side?

OpenStudy (superhelp101):

wait how did you get that. may i ask you to please explain?

OpenStudy (anonymous):

Expanding the binomial gives you \[(x+3)^2=x^2+6x+9\] The trick to completing the square is to find a number (in this case 9) such that when you take the square root and double it, you get the coefficient of the \(x\) term.

OpenStudy (anonymous):

In reverse order, you're given the coefficient, so you would cut it in half and square it: \[\left(\frac{6}{2}\right)^2=3^2=9\]

OpenStudy (superhelp101):

how I see you square it and got 9. Option 1

OpenStudy (superhelp101):

am I right?

OpenStudy (anonymous):

Yeah

OpenStudy (superhelp101):

Thank you for your help! :)